Monica flew home from vacation with a heavy bag. With the first airline she flew, Monica had to pay $17 to check her bag, plus $17 for every kilogram that her bag was over the weight limit. The next flight was with another airline that had the same weight limit. Monica had to pay $19 per kilogram that her bag was over the weight limit, in addition to the checked bag fee of $15. By coincidence, the fees ended up being the same with both airlines. How far over the weight limit was the bag? How large was each airline's fee?

Answer by ankor@dixie-net.com(22740)

Monica flew home from vacation with a heavy bag.
With the first airline she flew, Monica had to pay $17 to check her bag, plus $17 for every kilogram that her bag was over the weight limit.
The next flight was with another airline that had the same weight limit.
Monica had to pay $19 per kilogram that her bag was over the weight limit, in addition to the checked bag fee of $15.
By coincidence, the fees ended up being the same with both airlines.
How far over the weight limit was the bag?
How large was each airline's fee?
:
let x = the amt over the weight limit
17x + 17 = 19x + 15
17 - 15 = 19x - 17x
2 = 2x
x = 1 kg over the limit
"How large was each airline's fee?"
17(1) + 17 = $34
19(1) + 15 = $34